# C-DOG: Training-Free Multi-View Multi-Object Association in Dense Scenes Without Visual Feature via Connected $\delta$-Overlap Graphs thanks: Code available at GitHub: https://github.com/Yunghong/C-DOG

### v1

@@ IV-A Problem Formulation
The quality of the grouping shall be evaluated using the averaged square back projection error:
$E_{BP}=\sum_{\ell=1}^{L}\frac{1}{|G_{\ell}|}\sum_{(m,i_{m})\in G_{\ell}}\|V_{m,i_{m}}-\hat{V}_{m,i_{m}}\|^{2}$ (9)
where $V_{m,i_{m}}$ is the observed 2D feature point, and $\hat{V}_{m,i_{m}}$ is the back-projected 2D feature point.
@@
then one may conclude that $V_{m,i_{m}}$ and $V_{m^{\prime},i^{*}_{m^{\prime}}}$ are a pair of feature points associated with the same 3D point. Hence both of them should be assigned to the same association group. By repeating this process for every 2D points in every camera view, a desired grouping for the Point Feature Matching Problem may be obtained.
We remark that in rare cases, there may be more than one points on the $m^{\prime th}$ view satisfying eq. (11). This happens when the 3D points corresponding to these $V_{m^{\prime},i^{*}_{m^{\prime}}}$ 2D points fall on the same ray originated from the camera center of the $m^{th}$ camera view. The same $V_{m,i_{m}}$ then may be assigned to the corresponding association groups. We also consider the presence of outliers which are 2D feature points that are not included in any association groups. Each outlier corresponds to a 3D point which appears in only one camera view and hence stereo vision triangulation cannot be applied to estimate its 3D coordinate.
@@ IV-B C-DOG Overview
C-DOG is a training-free, geometry-aware association framework designed to serve as an intermediate module between 2D object detection (or pose estimation) and 3D reconstruction. In C-DOG, each 2D detection in a camera view is represented as a node in a graph. An edge will be established between a pair of nodes if these nodes are likely associated with the same 3D point. As illustrated in Fig. 4, the C-DOG consists of four main stages:
@@ IV-E Group-level Outlier Removal
IV-F Synthetic Benchmark Generation
@@ V Experiments
# views $\mathrm{G}$-$F_{1}$ $\mathrm{G}$-$\mathrm{IoU}$ $\mathrm{mP}$-$P$ $\mathrm{mP}$-$R$ $\mathrm{mP}$-$F_{1}$ $\mathrm{mP}$-$\mathrm{IoU}$ $\mathrm{PG}$-$P$ $\mathrm{PG}$-$R$ $\mathrm{PG}$-$F_{1}$ $3DErr$ $bpe$ time(ms)
@@ V-B Ablation Study
Method $\mathrm{G}$-$F_{1}$ $\mathrm{G}$-$\mathrm{IoU}$ $\mathrm{mP}$-$P$ $\mathrm{mP}$-$R$ $\mathrm{mP}$-$F_{1}$ $\mathrm{mP}$-$\mathrm{IoU}$ $\mathrm{PG}$-$P$ $\mathrm{PG}$-$R$ $\mathrm{PG}$-$F_{1}$ $3DErr$ $bpe$ time(ms)
@@
Method $\mathrm{G}$-$F_{1}$ $\mathrm{G}$-$\mathrm{IoU}$ $\mathrm{mP}$-$P$ $\mathrm{mP}$-$R$ $\mathrm{mP}$-$F_{1}$ $\mathrm{mP}$-$\mathrm{IoU}$ $\mathrm{PG}$-$P$ $\mathrm{PG}$-$R$ $\mathrm{PG}$-$F_{1}$ $3DErr$ $bpe$ time(ms)
@@
Method $\mathrm{G}$-$F_{1}$ $\mathrm{G}$-$\mathrm{IoU}$ $\mathrm{mP}$-$P$ $\mathrm{mP}$-$R$ $\mathrm{mP}$-$F_{1}$ $\mathrm{mP}$-$\mathrm{IoU}$ $\mathrm{PG}$-$P$ $\mathrm{PG}$-$R$ $\mathrm{PG}$-$F_{1}$ $3DErr$ $bpe$ time(ms)
@@
Method $\mathrm{G}$-$F_{1}$ $\mathrm{G}$-$\mathrm{IoU}$ $\mathrm{mP}$-$P$ $\mathrm{mP}$-$R$ $\mathrm{mP}$-$F_{1}$ $\mathrm{mP}$-$\mathrm{IoU}$ $\mathrm{PG}$-$P$ $\mathrm{PG}$-$R$ $\mathrm{PG}$-$F_{1}$ $3DErr$ $bpe$ time(ms)
@@
At the 10-camera setting, the method achieves the best performance in terms of group association, with a $5.4\%$ improvement in $\mathrm{G}$-$F_{1}$ compared to 2-view setups. In terms of runtime, the computational cost increases polynomially with the number of views, since more point pairs must be evaluated for mutual affinity.
@@ V-C Comparison With Existing Works
We selected methods that operate purely on spatial geometry, as discussed in Section II-B. These methods take the camera poses and 2D point coordinates in the image plane as input and produce association results in the form of grouped 2D points. The compared baselines include: the Greedy method [28], Composition-based Affinity Optimization (CAO) [36], ST-cut using 3D back-projection constraints (ST-Cut 3D bpj) [33], ST-cut using epipolar constraints (ST-Cut Epipolar) [32], bundle adjustment with soft matching (Bdl.Adj. w/ Sft. Mat.) [37], Connected Component Analysis (CCA) [40], Spectral Clustering [18], Permutation Synchronization (Permutation Sync.) [15], and Factorized Graph Matching (Fact. Graph Mat.) [16]. We also reported results for our proposed method under three different configurations: the original full model, and a variant without IQR-based outlier removal.
@@
Under noise-free conditions ($\sigma=0.00$), our method achieves a $13.3\%$ improvement over the second-best method in perfect group association ($PG$-$F_{1}$), a $2.8\%$ improvement in imperfect group association ($G$-$F_{1}$), and a $17.7\%$ gain in point-wise association performance ($mP$-$F_{1}$). In terms of $\mathrm{IoU}$, our method outperforms the second-best by $5.1\%$ for imperfect groups ($G$-$\mathrm{IoU}$) and $17.5\%$ for point-wise association quality ($mp$-$\mathrm{IoU}$). Under $\sigma=1.00$ conditions, our method achieves a $10.5\%$ improvement over the second-best method in $PG$-$F_{1}$, a $16.6\%$ gain in $mP$-$F_{1}$, a $0.1\%$ gain in $mP$-$F_{1}$, an $0.3\%$ gain in ($G$-$\mathrm{IoU}$), and an $16.8\%$ gain in ($mP$-$\mathrm{IoU}$). Under $\sigma=3.00$ conditions, our method achieves a $12.7\%$ improvement over the second-best method in $PG$-$F_{1}$, a $7.5\%$ gain in $mP$-$F_{1}$, and an $11.1\%$ gain in ($mP$-$\mathrm{IoU}$).
@@ VI Discussion
A key strength of our method is its robustness without relying on visual features, making it particularly well-suited for texture-poor or featureless scenarios. In addition, its training-free nature makes it ideal as an intermediate module that bridges upstream and downstream tasks—for example, serving as a correspondence associator between multi-view object detection and 3D reconstruction.
@@ VII Conclusion
We proposed an efficient, training-free method, C-DOG, for multi-view, multi-object association using only 2D point observations and known camera poses. Designed to operate effectively in featureless and noisy environments, our approach eliminates the need for visual features while preserving both scalability and accuracy. Extensive evaluations across varying noise levels, object densities, and view configurations demonstrate that our method consistently outperforms geometry-based baselines, achieving a strong balance between precision and runtime. This provides a general and robust solution that can serve as a bridging module between multi-view detection and 3D reconstruction. Future work will focus on improving computational efficiency and extending the approach to real-time applications in instance-rich environments.

### v2

@@ IV-A Problem Formulation
The quality of the grouping shall be evaluated using the averaged square back projection error defined in eq. (9) and eq. (10).
@@
then one may conclude that $V_{m,i_{m}}$ and $V_{m^{\prime},i^{*}_{m^{\prime}}}$ are a pair of feature points associated with the same 3D point. Hence both of them should be assigned to the same association group. By repeating this process for every 2D points in every camera view, a desired grouping for the Point Feature Matching Problem may be obtained. From eq. (5) we note that when two or more 3D points are colinear with respect to the camera center of the $m^{th}$ view, then we would also have
$d_{m,i_{m},m^{\prime},i^{\prime*}_{m^{\prime}}}=0$ (13)
We define outliers as 2D feature points that are not included in any association groups. An outlier may correspond to a 3D point which appears in only one camera view or an extremely noisy 2D feature point. An outlier 2D feature point will not be used for triangulation.
@@ IV-B C-DOG Overview
C-DOG is a geometry-aware association framework designed to serve as an intermediate module between 2D object detection (or pose estimation) and 3D reconstruction. In C-DOG, each 2D detection in a camera view is represented as a node in a graph. An edge will be established between a pair of nodes if these nodes are likely associated with the same 3D point. As illustrated in Fig. 4, the C-DOG consists of four main stages:
@@ IV-E Group-level Outlier Removal
IV-F Error Association Group Removal
While the previous step effectively removes intra-group outliers, some erroneous associations may still remain. Because the noise level of the algorithm is unknown, the acceptable BPE range cannot be predefined. To address this, we evaluate each group’s average BPE score, as defined in Equation 10, and remove entire groups when necessary to avoid incorporating erroneous reconstructions. Specifically, we first compute the BPE for each group and sort these scores in ascending order. We then examine consecutive groups, and if the difference between two successive BPE values exhibits a sudden increase, we discard the remaining groups. This approach provides an adaptive mechanism for identifying and removing erroneous association groups under unknown noise conditions.
IV-G Synthetic Benchmark Generation
@@ V Experiments
[…]
@@ V-B Ablation Study
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ V-C Comparison With Existing Works
We selected methods that operate purely on spatial geometry, as discussed in Section II-C. These methods take the camera poses and 2D point coordinates in the image plane as input and produce association results in the form of grouped 2D points. The compared baselines include: the Greedy method [43], Composition-based Affinity Optimization (CAO) [49], ST-cut using 3D back-projection constraints (ST-Cut 3D bpj) [46], ST-cut using epipolar constraints (ST-Cut Epipolar) [45], bundle adjustment with soft matching (Bdl.Adj. w/ Sft. Mat.) [50], Connected Component Analysis (CCA) [53], Spectral Clustering [24], Permutation Synchronization (Permutation Sync.) [21], and Factorized Graph Matching (Fact. Graph Mat.) [22]. We also reported results for our proposed method under three different configurations: the original full model, and a variant without IQR-based outlier removal.
@@
[…]
@@ VI Discussion
A key strength of our method is its robustness without relying on visual features, making it particularly well-suited for texture-poor or featureless scenarios. In addition, this makes it ideal as an intermediate module that bridges upstream and downstream tasks—for example, serving as a correspondence associator between multi-view object detection and 3D reconstruction.
@@ VII Conclusion
We proposed an efficient method, C-DOG, for multi-view, multi-object association using only 2D point observations and known camera poses. Designed to operate effectively in featureless and noisy environments, our approach eliminates the need for visual features while preserving both scalability and accuracy. Extensive evaluations across varying noise levels, object densities, and view configurations demonstrate that our method consistently outperforms geometry-based baselines, achieving a strong balance between precision and runtime. This provides a general and robust solution that can serve as a bridging module between multi-view detection and 3D reconstruction. Future work will focus on improving computational efficiency and extending the approach to real-time applications in instance-rich environments.